### v3

@@
Chenyu Huang1, Zhengyang Tang2,3, Dongdong Ge4, Shixi Hu5
@@ 6 Columbia University
rj2556@columbia.edu
{wangbenyou,wangzizhuo}@cuhk.edu.cn xin.zheng@duke.edu
@@
Optimization modeling and solving play a critical role in the application of Operations Research (OR) tools to address real-world problems, yet they pose challenges and require extensive expertise from OR experts. With the advent of large language models (LLMs), new opportunities have emerged to streamline and automate these tasks. However, current research predominantly relies on closed-source LLMs such as GPT-4, along with extensive prompt engineering techniques. This reliance stems from the scarcity of high-quality training datasets for optimization modeling, resulting in elevated costs, prolonged processing times, and privacy concerns. To address these challenges, our work is the first to propose a viable path for training open-source LLMs that are capable of optimization modeling as well as developing and executing solver codes, eventually leading to a superior ability for automating optimization modeling and solving. Particularly, we introduce a semi-automated data synthesis framework designed for optimization modeling issues, named OR-Instruct. This framework merges the training data requirements of large models with the unique characteristics of optimization modeling problems, and allows for customizable enhancements tailored to specific scenarios or modeling types. To evaluate the performance of our proposed framework, we present the IndustryOR benchmark, the inaugural industrial standard for evaluating LLMs in solving practical OR problems. Utilizing data synthesized through OR-Instruct, we train various open-source LLMs with a capacity of 7 billion parameters (dubbed ORLMs). The resulting model demonstrates significantly enhanced optimization modeling capabilities, achieving state-of-the-art performance across the NL4OPT, MAMO, and IndustryOR benchmarks. Our code and data are available at https://github.com/Cardinal-Operations/ORLM.
@@ 1 Introduction
The advent of large language models (LLMs) gives an emerging hope to the above challenge. As soon as its advent, these pre-trained LLMs have exhibited remarkable capabilities in rapid knowledge comprehension and cross-disciplinary learning, demonstrating remarkable success and transformative potential across various fields (Huang et al. 2024a). For instance, in the medical sector, the Hua Tuo large model significantly aids in diagnostics and treatment processes (Wang et al. 2023b). In the realm of software development, Codex (Zaremba et al. 2021) excels in numerous programming languages, specifically designed to enhance coding tasks. Additionally, these models are effectively utilized in automated perceptual analysis (Li et al. 2024), providing legal advice (Cui et al. 2023), offering stock market trading strategies (Chen et al. 2022b), and improving recommendation systems (Xu et al. 2024a). In more specialized scientific areas, such as mathematical derivations, LLMs also show substantial potential for advanced reasoning. For example, Xu et al. (2024b) develop a pipeline that incorporates self-critique mechanisms to refine the capabilities of LLMs in mathematical derivation. Furthermore, Trinh et al. (2024) train Alphaproof with hundreds of billions of synthetic data used to tackle complex olympiad geometry problems. Similarly, MathScale (Tang et al. 2024) offers a straightforward and scalable method to generate high-quality data for mathematical reasoning, establishing a foundation for further training of large language models in mathematical reasoning capabilities. Xing et al. (2024) and Ahmed and Choudhury (2024) both emphasize the specific challenges in translating linguistic descriptions into mathematical formulations. Their studies highlight the superiority of advanced LLMs like GPT-4 in handling complex optimization problems and the potential need for specialized datasets and finetuning approaches to further enhance performance.
@@
In the field of optimization modeling, a few studies have emerged in recent years attempting to automate the modeling and solving of optimization problems under natural language representations using pre-trained language models (PLMs) (Liu et al. 2019, Lewis et al. 2019). For example, NL4OPT (Ramamonjison et al. 2023) decomposes this task into multiple steps, where semantic entities are first recognized upon which mathematical models are formulated subsequently. However, this approach often results in accumulated errors in the pipeline and has limited generalization capabilities due to the relatively-small parameter scale. As LLMs become more advanced like the GPT-4 and Claude-3.5, researchers have begun to generate mathematical models by directly prompting these models. To extend to a complete solution that includes a program using mature solvers, prompting engineering methods such as Chain-of-Experts (Xiao et al. 2023), OptiMUS (AhmadiTeshnizi et al. 2024) and OptiGuide (Li et al. 2023a) employ multi-agent collaboration based on proprietary LLMs. These agents collaborate using complex reasoning chains to refine both the mathematical models and programs. Nonetheless, utilizing online APIs for advanced large models offers no assurances of privacy protection, which is prohibitive in some sensitive industry applications. Moreover, these approaches rely heavily on the capabilities of the foundational LLMs itself. As illustrated in Figure 1, even a basic optimization modeling problem presents challenges for the state-of-the-art foundational large language model, GPT-4. It demonstrates deficiencies in comprehending and transforming logical constraints, which imposes a limitation on methods dependent on these general closed-source LLMs.
@@
    To ensure the effectiveness, robustness, and real-world applicability of our model, we have identified four crucial requirements that the training dataset must fulfill, as informed by academic research like Alzubaidi et al. (2023), Jiang et al. (2024) and industry insights. Inspired by these requirements, we design and implement OR-Instruct, a semi-automated process for creating synthetic data tailored to optimization modeling. The process uses an iterative bootstrapping algorithm (see Figure 2). Initially, we collect a set of seed industry cases (e.g., 686 cases in our study) and add them to the training data pool. Following this, we use two strategies. One is expansion, which employs GPT-4 to generate data covering a wider range of scenarios and question types, expanding to thousands of scenarios and tens of question types. While this strategy broadens the coverage of scenarios, it does not ensure diversity in problem difficulty. GPT-4 struggles with independently solving complex problems and tends to simplify the optimization model, leading the strategy to generate results predominantly for easy and medium-level problems. To ensure that synthesized data retains the difficulty of the original seed data, we introduce the augmentation strategy. This strategy involves modifying objectives and constraints, rephrasing questions, and incorporating diverse modeling techniques. Essentially, we decompose the remaining requirements into subtasks, enabling GPT-4 to generate examples with greater problem-solution diversity within its capabilities. Finally, heuristics are used to automatically filter out obviously low-quality data. This cycle can be repeated multiple iterations until the training data pool reaches the requisite volume (e.g., 32,481 cases in our study). Importantly, this process is designed to facilitate customization by simply introducing new seed data and scenarios as needed.
@@
    We utilize the synthetic data generated by OR-Instruct to train a series of open-source large language models, each approximately 7 billion parameters in size, including Mistral-7B (Jiang et al. 2023), Deepseek-Math-7B-Base (Shao et al. 2024), and LLaMA-3-8B (AI@Meta 2024). We designate these trained models as Operation Research Language Models (ORLMs) and observe a marked enhancement in their optimization modeling capabilities. Our experimental results show that our best-performing ORLM achieves state-of-the-art performance on the NL4OPT, MAMO, and IndustryOR benchmarks.
To our best knowledge, our work is the first initiative to train open-source LLMs specifically for optimization modeling in real-world industries. This approach significantly reduces the reliance on closed-source LLMs thus adding significant flexibility to the training process and greatly enhancing privacy preservation for industrial applications. Remarkably, our highest-performing ORLM achieves state-of-the-art results on the NL4OPT, MAMO, and IndustryOR benchmarks. This indicates that even models with a parameter count as moderate as 7 billion can surpass the capabilities of larger models such as GPT-4 and comparable agent-based frameworks, following our specialized training methodology. This underscores the effectiveness of our approach.
We also want to comment that our proposed training method, the OR-Instruct, has the ability to facilitate customization to address particular needs of larger models, such as supply chain management, scheduling, inventory issues, or enhancing capabilities in linear programming, integer programming, and mixed integer programming. This synthetic data framework robustly addresses the scarcity of training data for optimization modeling, laying a solid foundation for future advancements in deep learning model training in this field.
@@
The structure of this paper is organized as follows: Section 2 reviews the literature pertinent to our research. Section 3 delineates four essential conditions required for the training data and outlines the implementation specifics of the semi-automated synthetic data framework for optimized modeling, OR-Instruct. Section 4 examines the impact of training open-source large models on various test sets and presents the experimental outcomes of customized enhancements in modeling capabilities. Section 5 compares the modeling effectiveness of GPT-4 with our ORLM and validates the necessity of the four key conditions through ablation studies. We conclude the paper in Section 6.
@@ 2 Literature Review
AI approach for facilitating solution to operations research problems. With the fast development of AI techniques, there have been a growing amount of research leveraging advanced techniques, such as deep learning, to address challenging operational research (OR) problems traditionally resistant to conventional methods. For example, in the supply chain context, Gijsbrechts et al. (2022) employ deep reinforcement learning to enhance inventory management, demonstrating parity with leading heuristics and approximate dynamic programming approaches. Han et al. (2023) examine a data-driven newsvendor problem resolved through deep neural networks (DNNs), offering theoretical assurances on excess risk bounds and validating the method’s effectiveness, Qi et al. (2023) investigate a data-driven multiperiod inventory replenishment problem with uncertain demand and vendor lead time, propose a one-step end-to-end framework that uses deep learning models to output the suggested replenishment amount directly from input features. In the revenue management field, Aouad and Désir (2022) propose a neural network-based discrete choice model called RUMnets, which combines the theoretical robustness of traditional economic models with the flexibility and learning capabilities of modern artificial intelligence approaches, leading to better predictions and insights into consumer behavior. Wang et al. (2023a) explore the deployment of deep neural networks in making the optimal assortment selection, demonstrating its effectiveness in capturing consumer choice. As for queuing theory, Dai and Gluzman (2022) investigate the deployment of advanced policy gradient methods for controlling multi-class queueing networks, underscoring the superior efficacy of the PPO algorithm under diverse load scenarios. For personalized recommendation, Wang et al. (2023c) utilize a novel unsupervised deep learning network and graph embedding techniques to improve recommendation accuracy and diversity by effectively transferring check-in patterns across different geographical contexts. Regarding general Markov decision process, Zhang et al. (2021) introduce a model-free reinforcement learning algorithm that attains asymptotically optimal sample complexity for learning an $\epsilon$-optimal policy within discounted Markov decision processes (MDPs). In addition, these techniques have been used in demand prediction, for example, Lee et al. (2024) propose a Transformer-based methodology for time series conformal prediction, utilizing its decoder to estimate prediction intervals through the quantile prediction of residuals, showcasing enhanced performance over existing conformal prediction techniques. Chen et al. (2020) introduce a novel architecture of convolutional neural networks (CNNs) that significantly improves forecasting by modeling the conditional distribution of future series data directly, rather than relying on traditional autoregressive methods. AI technology also demonstrates significant strengths in facilitating the implementation of operations research, for instance, Liu et al. (2023) employ reinforcement learning combined with spatial data to optimize the allocation of resources and strategize for disease outbreak responses. Parmentier (2022) introduces a novel paradigm that integrates machine learning with operations research to efficiently approximate complex industrial problems by transforming them into classic, more tractable optimization challenges.
@@
In addition, there have also been recent attempts in using AI methods to empower optimization algorithm design. For instance, deep learning or machine learning can be used to accelerate conventional optimization algorithms for tackling mixed-integer programming problems (e.g., Chen et al. 2023b, Nair et al. 2020). Numerous researchers also explore end-to-end optimization problem-solving strategies based on deep learning methods, motivated by the high structural similarity among many optimization problems, with differences primarily in data distribution (e.g., Maragno et al. 2023, Chen et al. 2022a, Khalil et al. 2017). Compared to this work, the above stream of works studies the integration of AI approach and traditional OR approaches in solving OR problems. In contrast, the main focus of the present work is in the facilitation and automation of the modeling process of the optimization model, which is different from the above stream.
@@
Automated optimization modeling is an emerging area within the field of Decision Intelligence, which aims to automate mathematical modeling and solving tasks to facilitate efficient decision-making. Recently, a plethora of studies have emerged, aiming to bridge the gap between natural language processing and mathematical optimization. One of the pioneering endeavors in this field is the NL4OPT competition, which focuses on formulating mathematical models for optimization problems (Ramamonjison et al. 2023). In particular, their work introduces a two-step framework using PLMs and offers a widely used benchmark for intelligence mathematical modeling. Expanding on this, Xiao et al. (2023) introduce Chain-of-Experts, a multi-agent LLM framework to build optimization model that significantly outperforms GPT-4 by incorporating domain-specific knowledge and cooperative reasoning strategies. AhmadiTeshnizi et al. (2024) propose OptiMUS, a robust LLM-based agent designed to formulate and solve MILP problems. This system not only develops mathematical models but also assists in writing and debugging solver code, showcasing a comprehensive solution that extends beyond simple problem formulation. In a similar vein, Li et al. (2023a) develop a framework called OptiGuide which leverages LLMs to interpret and explain optimization outcomes in supply chain management, thereby enhancing stakeholder comprehension and trust. In terms of diagnosing and resolving optimization issues, Chen et al. (2023a) introduce OptiChat, a natural language-based system equipped with GPT-4, which assists in identifying and addressing infeasible optimization models. This tool significantly aids both experts and novices in understanding and refining their models. Unlike these methods, our approach involves training open-source LLMs and delivering a complete solution using direct prompting, which is more concise and privacy-preserving than multi-agents interaction and simply calling API from advanced LLMs like GPT-4.
@@
Synthetic Data is one of the most important methods to improve the performance of large language models, with the advantage of low cost and diversity. A review of literature on synthetic data and LLMs is referred to Ding et al. (2024). In the realm of synthetic data for optimization modeling, traditional approaches often necessitate specialized expertise to transform real-world issues into mathematical formulations, thereby constraining broader accessibility and adaptability. Predominantly, researches explore how to automatically synthesize mixed-integer linear programming (MILP) to address these challenges effectively. For example, the methods proposed by Pawlak and Krawiec (2017) and Sroka and Pawlak (2018) utilize MILP examples for the synthesis of constraints. These methods develop an approach that can generate constraints from examples of both feasible and infeasible solutions, thereby demonstrating the feasibility of automating this conversion process. Further expanding on this theme, Pawlak and O’Neill (2021) leverage evolutionary strategies to generate constraints that improve the adaptability and efficiency of optimization models, thus enhancing the model’s flexibility in handling complex scenarios. Prasath and Karande (2023) examine the effectiveness of synthesizing mathematical programs from natural language specifications based on NL4OPT competition dataset, focusing on data synthesis methods using the CodeT5 model with enhancements from data augmentation and post-processing. The study highlights how synthetic data generation, notably through techniques like reverse translation and parameter variation, significantly improves model performance in solving optimization problems. The proposed synthetic data framework in Li et al. (2023b) leverages LLMs to identify decision variables, classify constraints, and generate MILP models based on NL4OPT problems. This approach significantly outperforms traditional methods, achieving higher accuracy in constraint recognition and model formulation. In contrast to the aforementioned synthetic data approaches, the method proposed in this paper is designed for the general purpose of synthesizing high quality data for optimization modeling from unstructured natural language descriptions. This allows for the development of synthesis strategies that are not only tailored to specific problem but also capable of customized data augmentation.

### v4

@@
Chenyu Huang1111* These authors contributed equally to this work., Zhengyang Tang2,3111* These authors contributed equally to this work., Shixi Hu5, Ruoqing Jiang6
@@ 6 Columbia University
xin.zheng@duke.edu ddge@sjtu.edu.cn
{wangbenyou,wangzizhuo}@cuhk.edu.cn
@@
Optimization modeling plays a critical role in the application of Operations Research (OR) tools to address real-world problems, yet they pose challenges and require extensive expertise from OR experts. With the advent of large language models (LLMs), new opportunities have emerged to streamline and automate such task. However, current research predominantly relies on closed-source LLMs such as GPT-4, along with extensive prompt engineering techniques. This reliance stems from the scarcity of high-quality training datasets for optimization modeling, resulting in elevated costs, prolonged processing times, and privacy concerns. To address these challenges, our work is the first to propose a viable path for training open-source LLMs that are capable of optimization modeling and developing solver codes, eventually leading to a superior ability for automating optimization modeling and solving. Particularly, we introduce OR-Instruct, a semi-automated data synthesis framework for optimization modeling that enables customizable enhancements for specific scenarios or model types. We also introduce IndustryOR, the first industrial benchmark for evaluating LLMs in solving practical OR problems. We train several 7B-scale open-source LLMs using synthesized data (dubbed ORLMs), which exhibit significantly enhanced optimization modeling capabilities, achieving state-of-the-art performance across the NL4OPT, MAMO, and IndustryOR benchmarks. Additionally, our experiments highlight the potential of scaling law and reinforcement learning to further enhance the performance of ORLMs. The workflows and human-machine interaction paradigms of ORLMs in practical industrial applications are also discussed in the paper.
@@ 1 Introduction
The advent of large language models (LLMs) gives an emerging hope to the above challenge. As soon as its advent, these pre-trained LLMs have exhibited remarkable capabilities in rapid knowledge comprehension and cross-disciplinary learning, demonstrating remarkable success and transformative potential across various fields (Huang et al. 2024a), like medical consultation (Wang et al. 2023b), software development (Zaremba et al. 2021), offering stock market trading strategies (Chen et al. 2022b), and improving recommendation systems (Xu et al. 2024). In the realm of more abstract mathematical reasoning, research has increasingly incorporated more techniques to enhance the capabilities of large models, yielding promising results. For example, Trinh et al. (2024) train Alphaproof with hundreds of billions of synthetic data used to tackle complex olympiad geometry problems. Similarly, MathScale (Tang et al. 2024) offers a straightforward and scalable method to generate high-quality data for mathematical reasoning, establishing a foundation for further training of large language models in mathematical reasoning capabilities.
@@
In recent years, with the advancements in pre-trained language models (PLMs) such as GPT-4 and Claude-3.5, a growing number of studies have explored automating the modeling and solving of optimization problems expressed in natural language. To produce complete solutions, including executable programs that leverage established solvers, existing approaches often employ prompt engineering techniques and multi-agent collaboration frameworks built on proprietary LLMs, such as Chain-of-Experts (Xiao et al. 2023), OptiMUS (AhmadiTeshnizi et al. 2024), and OptiGuide (Li et al. 2023a). However, utilizing online APIs for advanced large models offers no assurances of privacy protection, which is prohibitive in some sensitive industry applications. Additionally, these approaches are heavily reliant on the inherent capabilities of the foundational LLMs themselves, rendering them incapable of addressing the customized needs of users. As illustrated in Figure 1, even a basic optimization modeling problem presents challenges for the state-of-the-art foundational large language model, GPT-4. It demonstrates deficiencies in comprehending and transforming logical constraints, which imposes a limitation on methods dependent on these general closed-source LLMs.
@@
    To ensure the effectiveness, robustness, and real-world applicability of our model, we have identified four crucial requirements that the training dataset must fulfill, as informed by academic research like Alzubaidi et al. (2023), Jiang et al. (2024) and industry insights. Inspired by these requirements, we design and implement OR-Instruct, a semi-automated process for creating synthetic data tailored to optimization modeling. The process uses an iterative bootstrapping algorithm (see Figure 2). Initially, we collect a set of seed industry cases (e.g., 686 cases in our study) and add them to the training data pool. Following this, we use two strategies. One is expansion, which employs GPT-4 and leverages its in-context learning capability to generate data spanning diverse scenarios and question types. Another is the augmentation strategy, which involves modifying objectives and constraints, rephrasing questions, and incorporating diverse modeling techniques to ensure the difficulty and quality of generation problems. Finally, heuristics are used to automatically filter out obviously low-quality data. This cycle can be repeated multiple iterations until the training data pool reaches the requisite volume (e.g., 32,481 cases in our study). Notably, this approach is highly customizable; new seed data and scenarios can be seamlessly introduced to adapt the dataset to specific needs.
@@
    We utilize the synthetic data generated by OR-Instruct to train a series of open-source large language models, each approximately 7 billion parameters in size, including Mistral-7B (Jiang et al. 2023), Deepseek-Math-7B-Base (Shao et al. 2024), and LLaMA-3-8B (AI@Meta 2024). We designate these trained models as Operation Research Language Models (ORLMs) and observe a marked enhancement in their optimization modeling capabilities. Our experimental results show that our best-performing ORLM achieves state-of-the-art performance on the NL4OPT, MAMO, and IndustryOR benchmarks. Building upon this foundation, we conduct extensive numerical experiments to analyze the limitations of the ORLM and propose feasible solutions for its improvement. Our results highlight the promising prospects of ORLM in advancing the application and development of operations research, while also suggesting potential human-LLM collaboration frameworks. These frameworks demonstrate how ORLM could enhance the productivity of both industry and academia in the field of operations research.
To our best knowledge, our work is the first initiative to train open-source LLMs specifically for optimization modeling in real-world industries. This approach significantly reduces the reliance on closed-source LLMs thus adding significant flexibility to the training process and greatly enhancing privacy preservation for industrial applications. Remarkably, our highest-performing ORLM achieves state-of-the-art results on the NL4OPT, MAMO, and IndustryOR benchmarks. This suggests that in the field of optimization modeling, our approach enables ORLMs with approximately 7 billion parameters to outperform many cutting-edge large language models, such as GPT-4, Llama-3.1, Qwen 2.5, and related agent-based frameworks.
We also conduct a thorough analysis of the ORLM’s results, with numerous ablation experiments validating the effectiveness of our approach. Our findings show that, although ORLM performs relatively weak on some challenging datasets, several strategies can still ensure improvements in its performance. For instance, applying scaling laws to further increase the model size and data scale leads to continuous performance enhancements. Additionally, we find that ORLM has weak ranking capabilities for optimal optimization models. Using alignment strategies, such as reinforcement learning, also helps strengthen ORLM’s performance.
We also want to comment that our proposed data synthesis method, the OR-Instruct, has the ability to facilitate customization to address particular needs of larger models, such as supply chain management, scheduling, inventory issues, or enhancing capabilities in linear programming, integer programming, and mixed integer programming. This synthetic data framework robustly addresses the scarcity of training data for optimization modeling, provides a valuable reference for advancing the application of operations research in various industrial sectors with the support of large models. In Section 6.1, we provide a detailed discussion of representative workflows driven by large models..
@@
The structure of this paper is organized as follows: Section 2 reviews the literature related to our research. Section 3 delineates four essential conditions required for the training data and outlines the implementation specifics of the semi-automated synthetic data framework for optimized modeling, OR-Instruct. Section 4 examines the impact of training open-source large models on various test sets and presents the experimental outcomes of customized enhancements in modeling capabilities. Section 5 analyzes the limitations of ORLM through extensive experiments and discusses potential areas for improvement. Section 6 discusses the practical application scenarios of ORLM and explores future research directions based on the conclusions drawn in this paper. We conclude the paper in Section 7.
@@ 2 Literature Review
AI approach for facilitating solution to operations research problems. With the fast development of AI techniques, there have been a growing amount of research leveraging advanced techniques, such as deep learning, to address challenging operational research (OR) problems traditionally resistant to conventional methods. For example, in the supply chain context, Gijsbrechts et al. (2022) employ deep reinforcement learning to enhance inventory management, demonstrating parity with leading heuristics and approximate dynamic programming approaches. Qi et al. (2023) investigate a data-driven multi-period inventory replenishment problem with uncertain demand and vendor lead time, propose a one-step end-to-end framework that uses deep learning models to output the suggested replenishment amount directly from input features. In the revenue management field, Aouad and Désir (2022) propose a neural network-based discrete choice model called RUMnets, which combines the theoretical robustness of traditional economic models with the flexibility and learning capabilities of modern artificial intelligence approaches, leading to better predictions and insights into consumer behavior. Wang et al. (2023a) explore the deployment of deep neural networks in making the optimal assortment selection, demonstrating its effectiveness in capturing consumer choice. As for queuing theory, Dai and Gluzman (2022) investigate the deployment of advanced policy gradient methods for controlling multi-class queueing networks, underscoring the superior efficacy of the PPO algorithm under diverse load scenarios. For personalized recommendation, Wang et al. (2023c) utilize a novel unsupervised deep learning network and graph embedding techniques to improve recommendation accuracy and diversity by effectively transferring check-in patterns across different geographical contexts. Regarding general Markov decision process, Zhang et al. (2021) introduce a model-free reinforcement learning algorithm that attains asymptotically optimal sample complexity for learning an $\epsilon$-optimal policy within discounted Markov decision processes (MDPs). In addition, these techniques have been used in demand prediction, for example, Lee et al. (2024) propose a Transformer-based methodology for time series conformal prediction, utilizing its decoder to estimate prediction intervals through the quantile prediction of residuals, showcasing enhanced performance over existing conformal prediction techniques. Chen et al. (2020) introduce a novel architecture of convolutional neural networks (CNNs) that significantly improves forecasting by modeling the conditional distribution of future series data directly, rather than relying on traditional autoregressive methods.
AI technology also demonstrates significant strengths in facilitating the implementation of operations research, for instance, Liu et al. (2023) employ reinforcement learning combined with spatial data to optimize the allocation of resources and strategize for disease outbreak responses. Parmentier (2022) introduces a novel paradigm that integrates machine learning with operations research to efficiently approximate complex industrial problems by transforming them into classic, more tractable optimization challenges.
@@
In addition, there have also been recent attempts in using AI methods to empower optimization algorithm design. For instance, deep learning or machine learning can be used to accelerate conventional optimization algorithms for tackling mixed-integer programming problems (e.g., Chen et al. 2023c, Nair et al. 2020). Numerous researchers also explore end-to-end optimization problem-solving strategies based on deep learning methods, motivated by the high structural similarity among many optimization problems, with differences primarily in data distribution (e.g., Chen et al. 2022a, Khalil et al. 2017). Compared to this work, the above stream of works studies the integration of AI approach and traditional OR approaches in solving OR problems. In contrast, the main focus of the present work is in the facilitation and automation of the modeling process of the optimization model, which is different from the above stream.
@@
Automated optimization modeling is an emerging area, which aims to automate mathematical modeling and solving tasks to facilitate efficient decision-making. Recently, a growing body of studies have emerged, aiming to bridge the gap between natural language processing and mathematical optimization. One of the pioneering endeavors in this field is the NL4OPT competition, which focuses on formulating mathematical models for optimization problems (Ramamonjison et al. 2023). In particular, their work introduces a two-step framework using PLMs and offers a widely used benchmark for intelligence mathematical modeling. Expanding on this, Xiao et al. (2023) introduce Chain-of-Experts, a multi-agent LLM framework to build optimization model that significantly outperforms GPT-4 by incorporating domain-specific knowledge and cooperative reasoning strategies. AhmadiTeshnizi et al. (2024) propose OptiMUS, a robust LLM-based agent designed to formulate and solve MILP problems. This system not only develops mathematical models but also assists in writing and debugging solver code, showcasing a comprehensive solution that extends beyond simple problem formulation. Similarly, Li et al. (2023a) develop a framework called OptiGuide which leverages LLMs to interpret and explain optimization outcomes in supply chain management, thereby enhancing stakeholder comprehension and trust. In terms of diagnosing and resolving optimization issues, Chen et al. (2023a) introduce OptiChat, a natural language-based system equipped with GPT-4, which assists in identifying and addressing infeasible optimization models. Unlike these methods, our approach involves fine-tuning open-source LLMs and delivering a complete solution using direct prompting, which is more concise and privacy-preserving than multi-agents interaction and simply calling API from advanced LLMs like GPT-4.
@@
Synthetic data plays a crucial role in enhancing the capabilities of large language models by providing controlled, high-quality datasets that help fill gaps in real-world data, eliminate biases, and improve model robustness. A review of literature on synthetic data and LLMs is referred to Ding et al. (2024). In the field of synthetic data for optimization modeling, traditional methods often require specialized knowledge to convert real-world problems into mathematical models, which limits their accessibility and flexibility. Most research focuses on how to automatically generate mixed-integer linear programming (MILP) models to overcome these challenges more effectively. For example, the methods proposed by Pawlak and Krawiec (2017) and Sroka and Pawlak (2018) utilize MILP examples for the synthesis of constraints. These methods develop an approach that can generate constraints from examples of both feasible and infeasible solutions, thereby demonstrating the feasibility of automating this conversion process. Further expanding on this theme, Pawlak and O’Neill (2021) leverage evolutionary strategies to generate constraints that improve the adaptability and efficiency of optimization models, thus enhancing the model’s flexibility in handling complex scenarios. Prasath and Karande (2023) examine the effectiveness of synthesizing mathematical programs from natural language specifications based on NL4OPT competition dataset, focusing on data synthesis methods using the CodeT5 model with enhancements from data augmentation and post-processing. The study highlights how synthetic data generation, notably through techniques like reverse translation and parameter variation, significantly improves model performance in solving optimization problems. The synthetic data framework proposed in Li et al. (2023b) leverages LLMs to identify decision variables, classify constraints, and generate MILP models based on NL4OPT problems. This approach significantly outperforms traditional methods, achieving higher accuracy in constraint recognition and model formulation. In contrast to the aforementioned synthetic data approaches, the method proposed in this paper is designed for the general purpose of synthesizing high quality data for optimization modeling from unstructured natural language descriptions. This allows for the development of synthesis strategies that are not only tailored to specific problem but also capable of customized data augmentation.